40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 8 04:13:12 2012, kiwamu, Update, LSC, update on the locking activity 
    Reply  Fri Mar 9 13:34:13 2012, keiko, Update, LSC, update on the locking activity 
       Reply  Sat Mar 10 02:00:03 2012, keiko, Update, LSC, update on the locking activity 
Message ID: 6386     Entry time: Thu Mar 8 04:13:12 2012     Reply to this: 6393
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: update on the locking activity 

[Keiko / Kiwamu]

 Some updates on the locking activity:

  • Started summarizing the data of the Michelson lock in a wiki page:
  • Gradually moving on to the PRMI lock
    • The lock stays for reasonably a long time (~20 min or more)
    • POP22/110 demod signals seemed just ADC noise.
    • A first noise budget is in process
      • The glitches make the noise level worse above 40 Hz or so in both the MICH and PRCL budgets.
    • Sensing matrix will be measured tomorrow
    • The data will be also summarized in a wiki page
ELOG V3.1.3-